# A Survey on SAR ship classification using Deep Learning

### v1

@@ IV-B Handcrafted Features
Traditional features like histogram of oriented gradients (HOG) can be combined with features extracted by a CNN to enhance the performance of DL models that require a large amount of features. HOG features capture the distribution of edge orientations within an image region, describing its local shape and appearance. [31] uses a VGG network as the backbone for feature extraction and concatenate these features with HOG features on three categories (bulk carrier, container ship and tanker) from the OpenSARship dataset. The results indicate that the model that incorporates HOG features achieves 7% better performance than the model without HOG features.
@@
[43] proposes a novel method called HOG-ShipCLSNet. This method first extracts HOG features and then learns deep features using a CNN called the multiscale classification mechanism (MS-CLS). MS-CLS extracts features from each convolutional layer. Since features from different layers have varying dimensions, the network employs a global self-attention mechanism to focus on important features within each layer. To ensure stable training, the fully connected balance mechanism resizes features to a uniform size and then averages them. Finally, the HOG feature fusion mechanism merges the handcrafted HOG features with the DL features before performing classification. The authors conducted experiments on two datasets: three categories from OpenSARShip (bulk carrier, container ship, tanker) and seven categories from FUSARShip (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, other). The results indicate that HOG-ShipCLSNet is effective for ship classification. [67] proposes a novel attention-guided neural network for ship classification using the OpenSARship dataset. This network incorporates 14 handcrafted features processed by a feed-forward network and combines them with features extracted by the attention-guided network. These combined features are then employed for training a classification model. The model, trained on 4 categories from OpenSARship, is evaluated on a dataset of SAR-AIS integrated Sentinel-1 SAR images. This approach achieves an accuracy of 86.3%, demonstrating the potential effectiveness of combining handcrafted features with an attention mechanism for ship classification tasks.
@@
Choosing effective features plays a vital role in classification performance: meaningful features help the model to be trained optimally. [60] compares four different handcrafted features (Canny, Harris, Gabor, and local binary patterns (LBP)) and proposes a novel method for feature fusion that optimizes this process. The method consists of two blocks: (1) the knowledge supervision and collaboration module (KSCM), which facilitates mutual learning between handcrafted and deep features, and (2) the feature fusion and contribution assignment module (FFCS), which enhances channel attention. The proposed model has been tested on three categories from OpenSARShip and seven categories from FUSARShip using both deep and shallow versions of AlexNet, VGGNet, ResNet, and DenseNet. The results demonstrate that handcrafted features enhance the performance of all models, with Gabor features showing the most significant average feature gain of 3.4%. Additionally, shallow networks (VGG-11 and ResNet-19) exhibit the largest improvement over their baseline counterparts following feature fusion.
@@
[37] investigates the effect of combining handcrafted features with DL features by considering four handcrafted features (HOG, naive geometric features (NGF), local radar cross section (LRCS), and principal axis features (PAFs)) in conjunction with four DL models (AlexNet, VGGNet, ResNet, and DenseNet). They analyze different feature injection locations and techniques, recommending injection in the final fully-connected layer with dimension unification weighted concatenation and feature normalization (DUW-Cat-FN) for optimal results. The authors utilize three classes (bulk carrier, container ship, and tanker) from OpenSARShip and seven classes (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, and other) from FUSARShip. The results reveal that injecting handcrafted features into a DL model has the potential to improve performance by up to 6%, thus highlighting the continued importance of handcrafted features. [68] proposes a novel contrastive learning-based method that leverages the fusion of handcrafted features with SAR images. It employs data augmentation techniques on both SAR images, and handcrafted features, applying distinct augmentations to each data type. Subsequently, the model calculates three types of losses during deep training to facilitate learning of two primary pretext tasks and a cluster-level task. The proposed model, named DCPNet, achieves accuracies of 73.66% and 87.94% for 3-class and 7-class classifications on the OpenSARship and FUSAR-Ship datasets, respectively. These results suggest the potential effectiveness of contrastive learning for ship classification tasks.
@@
Images in OpenSARship contain information such as the length and breadth of the vessel. [32] enhances the ship classification performance of a DL model by combining the output of a CNN and a KNN (K-nearest neighbour classifies new data based on the majority class of its nearest neighbors). The KNN is trained to classify ships based on length and breadth taken from AIS data, while the CNN uses the image data. Three classes (cargo, tanker, and others) from OpenSARship dataset are used. The model combined with KNN showed superior performance compared to the CNN model alone.
@@
Since a ship comprises various sections, [41] first employees a minimum bounding rectangle (MBR) to detect the ship in the image. Then, subdivide the bounding box into the bow, midship, and stern sections and extract the features using two models based on the VGG network. One model leverages the data from the three sections to capture local features, while the other analyzes the entire MBR to capture global features. Finally, feature fusion is employed to classify the ships into two categories: military and civilian. The method, applied to high-resolution images captured by GaoFen-3, achieves an accuracy of 96%.
@@ IV-C Use of SAR properties
Maritime monitoring represents a unique challenge in case of moving ships. Indeed, motion can cause the ship to appear shifted within the SAR image with respect to its actual geographical position. This information can be proficiently exploited for classification purposes. [35] proposes a complex-valued CNN (CV-MotionNet) that considers both amplitude and phase information from single-look complex SAR images. The experiments were conducted on two datasets: a five-class simulated dataset and a three-class real-world GaoFen-3 dataset containing cargo ships, tankers, and other categories. The results demonstrate the model’s effectiveness in classifying SAR targets.
@@
Sidelobes, faint echoes in SAR images caused by the finite length of the radar pulse, can hinder the performance of DL models for ship classification. [38] investigates the effect of sidelobe elimination on ten classes (anti-pollution, bulk carrier, container, dredger, fishing, cargo, law enforce, passenger, oil tanker, and port tender) from FUSARShip using a simple CNN model. The results show that supressing the sidelobes improves the model performance, achieving up to 98% accuracy, which indicates the effectiveness of sidelobe elimination.
@@
[40] investigates whether using information from two different polarizations (VV and VH) can improve the model performance by capturing more features. The authors propose a novel network called squeeze-and-excitation Laplacian pyramid network with dual-polarization feature fusion (SE-LPN-DPFF). This network first merges features extracted from both polarizations. Then, it utilizes squeeze and excitation blocks to enhance feature representation by selectively emphasizing informative features. Finally, the features are fed into a Laplacian pyramid network based on residual blocks, which extracts information from different resolutions and performs classification ensuring that the model is trained on this fine-grained information. The authors employ two datasets from OpenSARship: one with three categories (bulk carriers, container ships, and tankers) and another with six categories (bulk carriers, cargos, container ships, fishing, general cargos and tankers). To ensure a balanced dataset, they use 70% of the data from the category with the fewest samples as the training set and the remaining data for testing. They compare the performance of SE-LPN-DPFF to 35 other DL classifiers. The results show that SE-LPN-DPFF outperforms all other models.
@@
[46] proposes a different approach using a dual-polarized network based on a pretrained VGG architecture. This network learns representations from both VV and VH polarizations and then apply bilinear pooling. They train the model on a three-class dataset (tanker, container ship, and bulk carrier) from OpenSARShip, where VV and VH images are paired for training. The results show that the model outperforms state-of-the-art models, achieving an accuracy of 88%.[49, 56] propose a novel methodology combining a loss function and bilinear pooling, termed the group bilinear convolutional neural network (GBCNN), based on dense blocks. Grouped VV and VH images are used to train two parallel models. The extracted features are then passed through bilinear pooling. The resulting features from both pooling layers are sent to the loss function, along with a third feature created by merging the pooling outputs. Utilizing three- and five-class (tanker, container ship, bulk carrier, cargo, general cargo) datasets from OpenSARShip with data augmentation techniques such as flipping, rotation, translation, and Gaussian noise, the GBCNN achieved accuracy scores of 88% and 67% for three and five classes, respectively.
@@
[51] investigates the polarization of images in conjunction with handcrafted features for ship classification. It proposes a novel polarization fusion-based geometric features network (PFGFE) that utilizes geometric features alongside VH-polarized, VV-polarized, and a combined VV-VH image. PFGFE employs three VGG network-based CNNs for feature extraction. A global feature self-attention mechanism is then applied to the output of each CNN. Subsequently, the extracted features are merged, and their classification probabilities are concatenated with corresponding coefficients. Geometric features are processed using LSTM to capture the importance of long-term dependencies. Finally, the outputs from the CNNs and LSTM are concatenated, and this final representation is used for classification. The authors evaluate PFGFE on three-class (bulk carrier, container ship, and tanker) and six-class (bulk carrier, cargo, container ship, fishing, general cargo, and tanker) datasets from OpenSARShip. The results demonstrate that PFGFE achieves accuracies of 79.84% and 56.83% respectively.
@@
Building on the idea that exploiting dual polarization information can enhance performance, [55] proposes a novel dual-polarization information-guided network (DPIG-Net) that leverages a polarization channel cross-attention framework (PCCAF) and a dilated residual dense learning framework (DRDLF) for effective feature extraction and fusion. It evaluates DPIG-Net on the OpenSARShip dataset using VH and VV polarizations for three (bulk carrier, container ship, tanker) and six classes (adding cargo, fishing, and general cargo), achieving accuracies of 81.28% and 58.68%, respectively. Smoothed polarization fusion, coupled with attention modules such as multiscale feature fusion and dynamic gating feature fusion, has been shown to improve model accuracy for ship classification. [77] employs this approach on OpenSARShip, achieving accuracies of 87.13% and 65.97% for three and six classes, respectively. These results suggest that smoothed polarization fusion may be more effective than coherence fusion [55]. [66] proposes a novel dual branch deep network for ship classification, tested on OpenSARship. The network leverages VH and VV polarizations and comprises three modules: (1) an image construction module that generates two pseudo-RGB images representing polarimetric and texture features, (2) a feature extraction module that employs two pre-trained ConvNeXt networks to extract features from the constructed images independently and (3) a feature fusion and classification module that combines the extracted features and performs classification. This network achieves accuracies of 87.62% and 64.26% for 3-class and 6-class classification tasks, respectively.
@@
The limited spatial resolution of SAR imagery presents a significant challenge for ship classification tasks, particularly for small vessels. These ships often occupy only a few pixels within the image, making it difficult to extract the features needed for an accurate classification. To address this, [26] proposes a joint convolutional network for enhancing the resolution of SAR images before classification. It uses TerraSAR-X images for pre-training and three classes (bulk carrier, container, tanker) from OpenSARship for training. The data are first fed into a generative model for upscaling, and then classified using a dense blocks-based convolutional network. It also compares the performance of models trained on VV and VH polarizations. The results show an increase in performance compared to simpler networks, achieving classification accuracies of 81.02% and 79.12% for VH and VV, respectively.
@@
When data are scarce and imbalanced, some DL methods, such as the siamese networks, can be helpful in overcoming this challenge. Siamese networks focus on learning a similarity function by training two networks that share weights, allowing for better learning of data representations. [28] proposes a novel loss function and apply a siamese network architecture. It utilizes two inception blocks on different polarizations (VV and VH) for four classes (cargo, tanker, tug, and other) from OpenSARship. The images are first passed through a target recognition algorithm (RCNN) to crop the targets. Then, the cropped images are fed into the inception blocks, the features are extracted and the loss is calculated. The results indicate that siamese networks with target detection outperform models without target detection, achieving an accuracy of 87.04%.
@@
In ship classification tasks, researchers typically use datasets with three or four categories. This is because DL models require a large amount of data for training, and maritime data are inherently imbalanced, making it challenging to achieve good performance with more classes. However, a study in [42] demonstrates the potential of feature fusion for handling a larger number of categories. It employees an eight-category dataset (cargo, tanker, carrier, container, fishing, dredger, tug, and passenger) from OpenSARShip. A VGG-based feature extractor is used to simultaneously extract feature maps from both VV and VH polarizations. These feature maps are then merged. Finally, a novel hybrid loss function was applied to reduce intra-class distances and increase inter-class distances, leading to improved classification performance. The results, with a mean average precision of 82%, support the effectiveness of this approach.
IV-D Fine-Tuning

### v2

@@ IV-B Handcrafted Features
Traditional features like histogram of oriented gradients (HOG) can be combined with features extracted by a CNN to enhance the performance of DL models that require a large amount of features. HOG features capture the distribution of edge orientations within an image region, describing its local shape and appearance. A VGG network is used in [31] as the backbone for feature extraction and concatenate these features with HOG features on three categories (bulk carrier, container ship and tanker) from the OpenSARship dataset. The results indicate that the model that incorporates HOG features achieves 7% better performance than the model without HOG features.
Handcrafted Feature Integration in SAR Ship Classification Feature Fusion (Handcrafted + DL)
HOG+CNN [31], HOG-ShipCLSNet [46],
Attention-guided handcrafted fusion [70],
Feature injection designs [37],
Mutual learning (KSCM+FFCS) [63]
Classical & Signal-Domain Features
Canny, Harris, Gabor, LBP [63],
Time–frequency SST $\rightarrow$ HOG + CNN [78]
Advanced Learning Strategies
Joint learning framework (FJL) [81],
Contrastive/self-supervised fusion (DCPNet) [71],
Bayesian CNNs for uncertainty [80]
Metadata & Structural Features
CNN+KNN with AIS length/breadth [32],
Part-based local+global features (MBR segmentation) [41]
Figure 4: Overview of handcrafted feature integration approaches in SAR ship classification. The methods are organized into four categories: (i) fusion of handcrafted descriptors with deep features, (ii) use of classical and signal-domain descriptors, (iii) advanced learning strategies incorporating handcrafted cues, and (iv) metadata- and structure-driven feature designs. Representative studies for each category are indicated by citation keys.
@@
A method called HOG-ShipCLSNet is proposed in [46]. This method first extracts HOG features and then learns deep features using a CNN called the multiscale classification mechanism (MS-CLS). MS-CLS extracts features from each convolutional layer. Since the features from different layers have varying dimensions, the network employs a global self-attention mechanism to focus on important features within each layer. To ensure stable training, a fully connected balance mechanism resizes the features to a uniform size and then averages them. Finally, the HOG feature fusion mechanism merges the handcrafted HOG features with the DL features before performing the classification. The authors conducted experiments on two datasets: three categories from OpenSARShip (bulk carrier, container ship, tanker) and seven categories from FUSARShip (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, other). The results indicate that HOG-ShipCLSNet is effective for ship classification. Another attention-guided neural network, using OpenSARship, is proposed in [70]. This network incorporates 14 handcrafted features processed by a feedforward network and combines them with features extracted by the attention-guided network. These combined features are then employed to train a classification model. The model, trained on 4 categories from OpenSARship, is evaluated on a dataset of SAR-AIS integrated Sentinel-1 SAR images. This approach achieves an accuracy of 86.3%, demonstrating the potential effectiveness of combining handcrafted features with an attention mechanism for ship classification tasks.
@@
Choosing effective features plays a vital role in classification performance: indeed, meaningful features help the model to be trained optimally. Four handcrafted features (Canny, Harris, Gabor, and local binary patterns (LBP)) are compared in [63], where an optimal method for feature fusion is proposed. The method consists of two blocks: (1) a knowledge supervision and collaboration module (KSCM), which facilitates mutual learning between handcrafted and deep features, and (2) a feature fusion and contribution assignment module (FFCS), which enhances channel attention. The proposed model has been tested on three OpenSARShip categories and seven FUSARShip categories using both deep and shallow versions of AlexNet, VGGNet, ResNet and DenseNet. The results show that handcrafted features improve the performance of all models, with Gabor features showing the most significant average gain of features of 3.4%. Shallow networks (VGG-11 and ResNet-19) exhibit the largest improvement over their baseline counterparts following feature fusion.
A ship classification framework that incorporates non-stationary information of SAR targets into DL models is introduced in [78]. By this approach, the data undergo the so-called “second-order synchrosqueezing transform” (SST), which allows the ship characteristics to be represented in the time–frequency domain. Then, the HOG features are extracted, thus reducing the dimensionality and providing discriminative inputs. These features are then fused with CNN representations, with ResNet identified as the most effective backbone. Experiments on 3 classes of OpenSARShip and FUSAR-Ship show clear performance gains, with accuracies of 82% and 90%, respectively. This study highlights the value of non-stationary signal characteristics for improving robustness and generalization in SAR ship classification.
In [81], a multidimensional feature joint learning framework (FJL-Framework) is introduced, designed to overcome the limitations of simple feature concatenation. The framework consists of three modules: (1) a subspace-level feature selection model with geometric constraints (SGC-GA) to refine pattern features, (2) a global–local interaction model (GFIC) to enhance deep feature learning, and (3) a joint learning model (MFJL) to fuse both types of features. This approach consistently achieves accuracies of 79.0% and 87.6% for 3-class OpenSARShip and 7-class FuSARShip respectively. By selectively integrating geometric, pattern, and deep features, the FJL-Framework demonstrates the importance of conflict-aware joint learning for robust SAR target recognition.
In [80], the use of Bayesian CNNs (BCNNs) for SAR ship classification is investigated, comparing them to deterministic CNNs. Using three classes from OpenSARShip, the study shows that a Bayesian variant of LeNet-5 (BLeNet-5) improves accuracy by 1.6% over LeNet-5, nearly matching the performance of VGG-16 while requiring far less computation. Beyond accuracy, BCNNs provide calibrated uncertainty estimates, decomposing them into aleatoric and epistemic components. The results suggest that Bayesian models can deliver more reliable predictions and offer useful confidence measures, highlighting their potential for robust SAR ship classification in limited-data settings.
@@
The effect of combining handcrafted features with DL features is investigated in [37] by considering four handcrafted features (HOG, naive geometric features (NGF), local radar cross section (LRCS), and principal axis features (PAFs)) in conjunction with four DL models (AlexNet, VGGNet, ResNet, and DenseNet). The authors analyze different feature injection locations and techniques, recommending injection in the final fully-connected layer with dimension unification, weighted concatenation and feature normalization (DUW-Cat-FN) for optimal results. They use three classes (bulk carrier, container ship, and tanker) from OpenSARShip and seven classes (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, and other) from FUSARShip. The results reveal that injecting handcrafted features into a DL model has the potential to improve performance by up to 6%, thus highlighting the importance of handcrafted features. In [71], a novel contrastive learning-based method is proposed, leveraging the fusion of handcrafted features with SAR images. It employs data augmentation techniques on both SAR images, and handcrafted features, applying distinct augmentations to each data type. Subsequently, the model calculates three types of losses during deep training to facilitate learning of two primary pretext tasks and a cluster-level task. The proposed model, named DCPNet, achieves accuracies of 73.66% and 87.94% for 3-class and 7-class classifications on the OpenSARship and FUSAR-Ship datasets, respectively. These results suggest the potential effectiveness of contrastive learning for ship classification tasks.
@@
Images in OpenSARship contain information such as the length and breadth of the vessel. The ship classification performance of a DL model is enhanced in [32] by combining the outputs of a CNN and a K-nearest neighbor model (KNN), where the latter classifies new data based on the majority class of its nearest neighbors. KNN is trained to classify ships based on length and breadth taken from AIS data, while the CNN uses the image data. Three classes (cargo, tanker, and others) from OpenSARship dataset are used. The model combined with KNN showed superior performance compared to the CNN model alone.
@@
Since a ship comprises various sections, the authors of [41] first employ a minimum bounding rectangle (MBR) to locate the ship in the image, then subdivide the bounding box into the bow, midship, and stern sections and extract the features using two models based on the VGG network. One model leverages the data from the three sections to capture local features, while the other analyzes the entire MBR to capture global features. Finally, feature fusion is employed to classify the ships into two categories: military and civilian. The method, applied to high-resolution images captured by GaoFen-3, achieves an accuracy of 96%.
@@ IV-C Use of SAR properties
Maritime monitoring through SAR imagery represents a unique challenge in case of moving ships. Indeed, motion can cause the ship to appear shifted within the SAR image with respect to its actual geographical position. This information can be exploited proficiently for classification purposes. A complex-valued CNN (CV-MotionNet) that considers both amplitude and phase information from single-look complex SAR images is proposed in [35]. The experiments were conducted on two datasets: a five-class simulated dataset and a three-class real-world GaoFen-3 dataset containing cargo ships, tankers, and other categories. The results demonstrate the model’s effectiveness in classifying SAR targets.
Use of SAR Properties in Ship Classification Motion & Phase–Physics (Complex SAR)
Complex-valued CNN for motion cues [35];
Phase-aware network (FDC–TA–DSN) with complex SAR [85]
Polarimetric Fusion (VV/VH)
SE–LPN–DPFF dual-pol fusion [40]; VGG+bilinear dual-pol [49];
GBCNN w/ grouped bilinear pooling [52, 59];
DPIG-Net (PCCAF, DRDLF) [58]; smoothed pol. fusion [76];
Dual-branch ConvNeXt w/ pseudo-RGB pol./texture [69];
Polarimetric interaction (CPINet, MO-SE, FBC) [77];
Pol.+geometric handcrafted fusion (PFGFE) [54]
Scattering/Signal Conditioning
Multiscale global scattering association (MGSFA-Net) [79];
Sidelobe suppression (preprocessing) [38]
Resolution & Regime Handling
Super-resolution $\rightarrow$ classify (TSX pretrain, OpenSARShip) [26];
Siamese + detection; dual-pol inputs [28];
Large-class dual-pol fusion (8-class) [45]
Figure 5: Overview of methods that explicitly exploit SAR imaging properties for ship classification. Approaches are grouped into four categories: (i) motion and phase information from complex-valued data, (ii) polarimetric fusion of VV/VH channels, (iii) scattering- and signal-conditioning techniques, and (iv) resolution- and regime-specific strategies. Representative studies for each category are indicated by citation keys.
@@
Sidelobes, faint echoes in SAR images caused by the finite length of the radar pulse, can hinder the performance of DL models for ship classification. The effect of sidelobe suppression is investigated in [38] using a simple CNN model on ten classes from FUSARShip (anti-pollution, bulk carrier, container, dredger, fishing, cargo, law enforce, passenger, oil tanker, and port tender). The results show that suppressing the sidelobes improves the model performance, achieving up to 98% accuracy.
@@
The authors of [40] investigate whether using information from two different polarizations, VV and VH, can improve the model performance by capturing more useful features. They propose a novel network called squeeze-and-excitation Laplacian pyramid network with dual-polarization feature fusion (SE-LPN-DPFF). This first merges features extracted from both polarizations, then uses squeeze and excitation blocks to enhance feature representation by selectively emphasizing informative features. Finally, the features are fed into a Laplacian pyramid network based on residual blocks, which extracts information from different resolutions and performs classification ensuring that the model is trained on this fine-grained information. Two datasets from OpenSARship are used, one with three categories (bulk carriers, container ships, and tankers) and another with six categories (bulk carriers, cargo, container ships, fishing, general cargo and tankers). To ensure a balanced dataset, the authors use 70% of the data from the category with the fewest samples as the training set and the remaining data for testing. They compare the performance of SE-LPN-DPFF to 35 other DL classifiers. The results show that SE-LPN-DPFF outperforms all other models.
@@
A different approach is proposed in [49], using a dual-polarized network based on a pretrained VGG architecture. This network learns representations from both VV and VH polarizations and then applies bilinear pooling. The authors train the model on a three-class dataset from OpenSARShip (tanker, container ship, and bulk carrier), where VV and VH images are paired for training. The results show that the model outperforms state-of-the-art models, achieving an accuracy of 88%. A novel methodology, named the group bilinear convolutional neural network (GBCNN), is proposed in [52, 59], combining a loss function and bilinear pooling, based on dense blocks. Grouped VV and VH images are used to train two parallel models. The extracted features are then passed through bilinear pooling. The resulting features from both pooling layers are sent to the loss function, along with a third feature created by merging the pooling outputs. Using three- and five-class datasets from OpenSARShip (tanker, container ship, bulk carrier, cargo, general cargo), with data augmentation techniques such as flipping, rotation, translation, and Gaussian noise, the GBCNN achieved accuracy scores of 88% and 67% for three and five classes, respectively.
Recent work has begun to explore the use of complex-valued SAR data for ship classification, leveraging both amplitude and phase information rather than relying on intensity alone. The authors of [85] introduce FDC-TA-DSN, a deep SAR network that combines four-dimensional dynamic convolution (FDC) to suppress speckle noise with a triple-attention (TA) mechanism to filter sea clutter and emphasize discriminative features. Phase information is incorporated through joint time–frequency analysis and autoencoder-based feature learning, and fused with amplitude features to enhance classification. To support this effort, a new high-resolution dataset, ComplexSAR_Ship, was constructed from Gaofen-3 imagery, covering 17 ship categories. Experiments show that FDC-TA-DSN consistently outperforms standard CNNs and DSN baselines, confirming the value of phase information and complex-valued features for more robust ship classification.
@@
In [54], polarization is investigated in conjunction with handcrafted features. A novel polarization fusion-based geometric features network (PFGFE) is proposed, utilizing geometric features alongside VH-polarized, VV-polarized and a combined VV-VH image. PFGFE employs three VGG network-based CNNs for feature extraction. A global feature self-attention mechanism is then applied to the output of each CNN. Subsequently, the extracted features are merged, and their classification probabilities are concatenated with corresponding coefficients. Geometric features are processed using LSTM to capture the importance of long-term dependencies. Finally, the outputs from the CNNs and LSTM are concatenated, and this final representation is used for classification. The authors evaluate PFGFE on three-class and six-class datasets from OpenSARShip (respectively, bulk carrier, container ship, and tanker, and bulk carrier, cargo, container ship, fishing, general cargo, and tanker). The results show accuracies of 79.84% and 56.83% respectively.
@@
Building on the idea that exploiting dual polarization information can enhance performance, the authors of [58] propose a novel dual-polarization information-guided network (DPIG-Net) that leverages a polarization channel cross-attention framework (PCCAF) and a dilated residual dense learning framework (DRDLF) for effective feature extraction and fusion. It evaluates DPIG-Net on the OpenSARShip dataset using VH and VV polarizations for three (bulk carrier, container ship, tanker) and six classes (adding cargo, fishing, and general cargo), achieving accuracies of 81.28% and 58.68%, respectively. Smoothed polarization fusion, coupled with attention modules such as multiscale feature fusion and dynamic gating feature fusion, has been shown to improve model accuracy for ship classification. This approach is adopted in [76] on OpenSARShip, achieving accuracies of 87.13% and 65.97% for three and six classes, respectively. These results suggest that smoothed polarization fusion can be more effective than coherence fusion [58]. In [69], a novel dual branch deep network for ship classification is proposed and tested on OpenSARship. The network leverages VH and VV polarizations and comprises three modules: (1) an image construction module that generates two pseudo-RGB images representing polarimetric and texture features, (2) a feature extraction module that employs two pretrained ConvNeXt networks to extract features from the constructed images independently and (3) a feature fusion and classification module that combines the extracted features and performs classification. This network achieves accuracies of 87.62% and 64.26% for 3-class and 6-class classification tasks, respectively.
A Multiscale Global Scattering Feature Association Network (MGSFA-Net) is proposed in [79]. It integrates scattering features with deep features for SAR ship classification. After fine segmentation to isolate ship targets, scattering centers (SCs) are extracted and modeled as local graphs, which are then associated through the SCFA module and enhanced with a multiscale fusion module. In parallel, deep features are extracted using a multikernel module, and both representations are fused by weighted integration. Experiments on FUSAR-Ship and OpenSARShip demonstrate that MGSFA-Net achieves accuracies of 85.3% (three-class FuSARShip) and 85.7% (three-class OpenSARShip), outperforming other feature-fusion methods. The results show that incorporating scattering characteristics improves robustness and interpretability, particularly in few-shot scenarios.
A Cross-Polarimetric Interaction Network (CPINet) is proposed in [77] to exploit the complementary information in dual-polarized SAR images. The network employs a multiscale deep feature extraction framework and introduces a mixed-order squeeze–excitation (MO-SE) attention module, where first- and second-order statistics from one polarization guide the learning of the other. To further enhance feature fusion, factorized bilinear coding (FBC) is adopted, which reduces the parameter size while suppressing noise. An adaptive loss balancing strategy based on GradNorm is also applied to optimize multitask training. Evaluated on three- and five-class subsets of OpenSARShip, CPINet consistently outperforms both single-pol baselines and prior dual-pol methods, demonstrating the effectiveness of polarimetric feature interaction for robust ship classification.
@@
The limited spatial resolution of SAR imagery presents a significant challenge for ship classification tasks, particularly for small vessels. These ships often occupy only a few pixels within the image, making it difficult to extract the features needed for an accurate classification. To address this, the authors of [26] propose a joint convolutional network to enhance the spatial resolution before classification. It uses TerraSAR-X images for pretraining and three classes from OpenSARship for training (bulk carrier, container, tanker). The data are first fed into a generative model for upscaling, and then classified using a dense blocks-based convolutional network. It also compares the performance of models trained on VV and VH polarizations. The results show an increase in performance compared to simpler networks, achieving classification accuracies of 81.02% and 79.12% for VH and VV, respectively.
@@
Methods such as the siamese networks can be helpful in overcoming the challenge of scarce and imbalanced data. Siamese networks focus on learning a similarity function by training two networks that share weights, allowing for better learning of data representations. In [28], a novel loss function with a siamese network architecture are proposed. Two inception blocks are used on different polarizations (VV and VH) for four classes from OpenSARship (cargo, tanker, tug, and other). The images are first passed through a target recognition algorithm (RCNN) to crop the targets. Then, the cropped images are fed into the inception blocks, the features are extracted and the loss is calculated. The results indicate that siamese networks with target detection outperform models without target detection, achieving an accuracy of 87.04%.
@@
In ship classification tasks, researchers typically use datasets with three or four categories. This is because DL models require a large amount of data for training, and maritime data are inherently imbalanced, making it challenging to achieve good performance with more classes. However, a study in [45] demonstrates the potential of feature fusion for handling a larger number of categories. It employs an eight-category dataset from OpenSARShip (cargo, tanker, carrier, container, fishing, dredger, tug, and passenger). A VGG-based feature extractor is used to simultaneously extract feature maps from both VV and VH polarizations. These feature maps are then merged. Finally, a novel hybrid loss function is applied to reduce intra-class distances and increase inter-class distances, leading to improved classification performance. The results, with a mean average precision of 82%, support the effectiveness of this approach.
IV-D Transfer Learning and Domain Adaptation
Transfer Learning and Domain Adaptation in SAR Ship Classification Fine-Tuning Pre-trained Models
VGG-16 fine-tuning (f1 = 97.8%) [22];
DenseNet-121 + augmentation (98%) [24];
Impact of incident angle variation [25]
Domain Adaptation
DSAN++ (optical $\rightarrow$ SAR) [47];
Cross-dataset adaptation (OpenSARShip $\leftrightarrow$ FUSARShip) [73];
Domain-specific vs. non-domain-specific pretraining [55]
Data Augmentation & Generative Models
Augmentation + handcrafted features [39];
Conditional GANs for high-res SAR [48];
Generative domain adaptation [68];
Attention-Dense CycleGAN (OPT2SAR) [66]
Figure 6: Overview of transfer learning strategies for SAR ship classification. The approaches are grouped into three categories: (i) fine-tuning of pre-trained models, (ii) domain adaptation across modalities or datasets, and (iii) data augmentation and generative models to mitigate data scarcity. Representative studies are indicated by citation keys.